# Numerical validation of scaling laws for stratified turbulence

### v1

@@ Abstract
Recent theoretical progress using multiscale asymptotic analysis has revealed various possible regimes of stratified turbulence. Notably, buoyancy transport can either be dominated by advection or diffusion, depending on the effective Péclet number of the flow. Two types of asymptotic models have been proposed, which yield measurably different predictions for the characteristic vertical velocity and length scale of the turbulent eddies in both diffusive and non-diffusive regimes. The first, termed a ‘single-scale model’, is designed to describe flow structures having large horizontal and small vertical scales, while the second, termed a ‘multiscale model’, additionally incorporates flow features with small horizontal scales, and reduces to the single-scale model in their absence. By comparing predicted vertical velocity scaling laws with direct numerical simulation data, we show that the multiscale model correctly captures the properties of strongly stratified turbulence within spatiotemporally-intermittent turbulent patches. Meanwhile its single-scale reduction accurately describes the more orderly layer-like flow outside those patches.
@@ 1 Introduction
Owing to the associated enhanced rates of irreversible scalar mixing, stratified turbulence is a critical process in the Earth’s atmosphere and oceans, impacting both weather and climate, and in the interiors of stars and gaseous planets, affecting their long-term evolution. Assuming that the buoyancy of the fluid is controlled by a single scalar field, which could be temperature or the concentration of a single solute, the dimensionless Boussinesq equations governing the fluid motions are
@@
$\boldsymbol{\nabla}\cdot\boldsymbol{u}$ $=0,$ (1c)
@@
where $\boldsymbol{u}=(u,v,w)$ is the velocity field expressed in units of $U$ (where $U$ is a characteristic horizontal velocity of the large-scale flow), $t$ is the time variable in units of $L/U$ (where $L$ is a characteristic large horizontal scale of the flow), $p$ is the pressure fluctuation away from hydrostatic equilibrium in units of $\rho_{m}U^{2}$ (where $\rho_{m}$ is the mean density of the fluid), and $b$ is the deviation of the buoyancy field away from a linearly stratified background, expressed in units of $LN^{2}$ (where $N$ is the buoyancy frequency of the stable stratification). The flow is assumed to be driven by a non-dimensional divergence-free horizontal force ${\boldsymbol{F}}_{h}$, which only varies on large spatial scales and long time scales. The unit vector $\boldsymbol{e}_{z}$ points in the direction opposite to gravity.
@@
The usual dimensionless governing parameters of the flow emerge; namely, the Reynolds number $Re=UL/\nu$, the Péclet number $Pe=UL/\kappa$ and the Froude number $Fr=U/NL$, where $\nu$ is the kinematic viscosity, and $\kappa$ is the buoyancy diffusivity, while the Prandtl number, $Pr=\nu/\kappa=Pe/Re$ is a property of the fluid. Typically, $Pr\sim O(1)$ in air and water, but is very small in astrophysical fluids (of order $10^{-2}$ in degenerate plasmas and liquid metals, and much smaller in non-degenerate stellar plasmas, see Lignières,, 2020).
@@
As the stratification increases ($Fr\rightarrow 0$), vertical motions are increasingly suppressed and restricted to small characteristic vertical scales $l_{z}=O(\alpha L)$, where the emergent aspect-ratio $\alpha$ is an increasing function of $Fr$ but could also depend on $Re$ and $Pe$. In the limit $(\alpha,Fr)\rightarrow 0$, asymptotic analysis has successfully been used to derive reduced equations for stratified turbulence and to gain insight into its properties. Brethouwer et al., (2007), following Billant and Chomaz, (2001), proposed an asymptotic reduction in which the vertical coordinate $z$ is rescaled as $\zeta=z/\alpha$ (with $\alpha\ll 1$), while the horizontal coordinates ${\boldsymbol{x}}_{h}=(x,y)$ remain of order unity. Accordingly, all dependent variables $q$ are expressed as $q({\boldsymbol{x}}_{h},\zeta,t;\alpha)$. As in Klein, (2010), we refer to this type of model, which is designed to capture the essence of a scale-specific process, as a single-scale asymptotic (SSA) model. This vertical rescaling, when used in (1), reveals the importance of the emergent buoyancy Reynolds and Péclet numbers, defined as
@@
respectively. Brethouwer et al., (2007) showed that balancing the mass continuity equation in the limit $\alpha\rightarrow 0$ requires $w=O(\alpha)$. When $Re_{b}$ and $Pe_{b}$ are at least $O(1)$, dominant balance in the buoyancy equation implies $b=O(\alpha)$. Finally, the vertical component of the momentum equation reduces to hydrostatic equilibrium when $\alpha\rightarrow 0$, yielding $\alpha=Fr$, as first argued in the inviscid and non-diffusive case by Billant and Chomaz, (2001).
More recently, Shah et al., (2024) noted that in the limit of $Pr\ll 1$, it is possible to have a regime in which $Pe_{b}\ll 1\leq Re_{b}$. They demonstrated that in this case, the SSA model and corresponding asymptotic expansion reveal instead that $w=O(\alpha)$ and $b=O(\alpha Pe_{b})$, with $\alpha=(Fr^{2}/Pe)^{1/4}$ (see also Lignières,, 2020; Skoutnev,, 2023).
Crucial to the SSA theory is the notion that every component of the flow is strongly anisotropic, with large horizontal scales and a small vertical scale. In this model, therefore, the vertical fluid motions are primarily driven by the divergence of the horizontal flow, as illustrated schematically in figure 1. Chini et al., (2022), however, noted that the SSA theory ignores the possibility that isotropic motions with small horizontal scales may also exist and, in fact, are commonly seen in numerical simulations of stratified turbulence at sufficiently large $Re_{b}$ (cf. Maffioli and Davidson,, 2016; Cope et al.,, 2020; Garaud,, 2020). They proposed a new asymptotic reduction that explicitly incorporates two horizontal scales and two time scales, such that all dependent variables are expressed as
@@
and where the subscripts $s$ and $f$ are used to denote slow and fast scales, respectively. Again following Klein, (2010), we refer to the resulting reduced equations as a multiscale asymptotic (MSA) model. Chini et al., (2022) showed that these definitions imply that the large-scale motions remain strongly anisotropic with an aspect ratio $\alpha$, as in the SSA model, but can coexist with isotropic small-scale motions that evolve on the fast time scale $t_{f}$, and vary on the small vertical coordinate $\zeta$ and the ‘fast’ horizontal coordinate ${\boldsymbol{x}}_{f}$. These small-scale motions are self-consistently driven by an instability of the local vertical shear emergent from the larger-scale horizontal flow structures (see figure 1), and are gradually stabilized as the stratification increases at fixed $Re$. They essentially disappear beyond a certain threshold, at which point the MSA model naturally recovers the SSA model and its predicted scalings. For the sake of clarity, however, we refer in what follows to the SSA model and its scalings whenever small horizontal scales are ignored, and to the MSA model and its scalings whenever they are taken into account, even though the MSA model does in fact naturally cover both cases.
@@
In the asymptotic limit where $Re_{b}\geq O(1)$ and $Pe_{b}\geq O(1)$, Chini et al., (2022) found that $w=O(\alpha^{1/2})$, $b=O(\alpha)$ and $\alpha=Fr$ when $\alpha\rightarrow 0$. Their scaling prediction for $w$ thus deviates substantially from that of Brethouwer et al., (2007), but recovers that of Riley and Lindborg, (2012) albeit using different arguments (for details see Shah et al.,, 2024). That scaling has been tentatively validated by Maffioli and Davidson, (2016) in run-down direct numerical simulations (DNS) of stratified turbulence.
@@
Extending the MSA theory to the low $Pr$ case, Shah et al., (2024) recovered the results of Chini et al., (2022) when $Pe_{b}\geq\mathit{O}(1)$. They also found that $\alpha=Fr$ and $w=O(\alpha^{1/2})$ both continue to hold in an ‘intermediate’ regime where $O(\alpha)\leq Pe_{b}\ll 1$. However, when $Pe_{b}\ll\alpha$, a new fully diffusive regime emerges in which $w=O(\alpha^{1/2})$, $b=O(Pe_{b}\alpha^{1/2})$ and $\alpha=(Fr^{2}/Pe)^{1/3}$. As for the $Pr=\mathit{O}(1)$ scenario, the scaling predictions of the MSA theory differ substantially from those emerging from the low $Pe_{b}$ limit of the SSA theory (Shah et al.,, 2024) but recover them when small scales are absent. The various theories and their predicted scalings for $Re_{b}\geq O(1)$, with $Pe_{b}\geq O(\alpha)$ or $Pe_{b}\ll\alpha$, respectively, are summarized in figure 1.
@@
Therefore, an interesting question is whether evidence for these scaling laws can be found in DNS data. Recently, two series of DNS were presented by Cope et al., (2020) and Garaud, (2020), respectively, which solved equations (1) with $\boldsymbol{F}_{h}=\sin(y)\boldsymbol{e}_{x}$ (where $\boldsymbol{e}_{x}$ is a unit vector in the streamwise, i.e. $x$, direction; see §2 for further details). In their $Pe<1$ simulations (where by construction $Pe_{b}\ll\alpha$), Cope et al., (2020) found that the vertical length scale of the turbulent motions scales as $\alpha=(Fr^{2}/Pe)^{1/3}$, validating the predictions of Shah et al., (2024) in that limit. This scaling, however, was not as clearly evident in the high $Pe$ but low $Pe_{b}$ data of Garaud, (2020). One potential explanation is that $Re_{b}$ is relatively low in these simulations (which have $Pr=0.1$, so $Re_{b}=10Pe_{b}$), implying viscous effects are not necessarily negligible. In the limit of high $Pe_{b}$, Garaud, (2020) was unable to find evidence for the $\alpha=Fr$, $w=O(Fr)$ scaling of Brethouwer et al., (2007) and was unaware at the time of the scaling $w=O(Fr^{1/2})$ obtained by Chini et al., (2022), proposing instead on empirical grounds that $w\propto\alpha=Fr^{2/3}$ provides the best fit to the data. The apparent discrepancy between Garaud’s data and previous models therefore prompts us to analyze some new DNS results and to revisit the available data from Cope et al., (2020) and Garaud, (2020) in the light of the MSA models of stratified turbulence recently derived by Chini et al., (2022) and Shah et al., (2024).
@@ 2 Comparison of theory with DNS
In what follows, we extend and re-analyze the datasets presented in Cope et al., (2020) and Garaud, (2020). Both studies performed DNS of the set of non-dimensional equations (1) with $\boldsymbol{F}_{h}=\sin(y)\boldsymbol{e}_{x}$ in a triply-periodic domain of size $4\pi\times 2\pi\times 2\pi$ using the PADDI code (Traxler et al.,, 2011). We focus on their highest Reynolds number simulations, which were performed for $Re=600$. Note that in this paper, Cope et al., (2020) and Garaud, (2020), $Re$ is defined using the inverse wavenumber of the forcing, and thus is a factor of $2\pi$ smaller than that of simulations which use the box size as the unit of length instead. Cope et al., (2020) presented a range of DNS for $Pe\leq 0.1$ and low $Fr$ (using the parameter $B=Fr^{-2}$ to characterize the stratification). They also ran a few simulations in the asymptotically low $Pe$ regime (Lignières,, 1999), called the LPN regime hereafter, where the buoyancy equation is replaced by $w=Pe^{-1}\nabla^{2}b$. Garaud, (2020) presented DNS with $Re=600$, $Pe=60$ (i.e. $Pr=0.1$) and low $Fr$. Each simulation was integrated until a statistically-stationary state was reached, lasting at least 100 time units. In some cases, we had to further extend the original DNS from Cope et al., (2020) or Garaud, (2020) to have a sufficiently long stationary time series. The quantity $w_{rms}=\langle w^{2}\rangle_{t}^{1/2}$, where $\langle\cdot\rangle_{t}$ denotes a volume and time average, was then measured in that statistically stationary state using the extended data.
@@
To complement this dataset, we have run additional simulations at $Re=1000$, $Pe=100$. These DNS have twice the spatial resolution of those of Cope et al., (2020) and Garaud, (2020) and, thus, have only been integrated for up to 50 time units in the statistically stationary regime. In addition, the full fields are too large to be saved regularly, so we have saved two-dimensional slices through the data in the $(x,y)$, $(y,z)$ and $(x,z)$ planes. These simulations are only used for visualizations and to assess the influence of viscosity by comparison with the $Re=600$ results.
@@
We compare the $w_{rms}$ data and the various theories in the top row of figure 2. The left panel shows $w_{rms}$ as a function of $Fr^{-1}$, measured for the $Pe=60$, $Re=600$ runs (green symbols), and for the $Pe=100$, $Re=1000$ runs (orange symbols). Note that $Pr=0.1$ in both cases. We refer to these simulations as ‘non-diffusive’ runs because $Pe$ is large. The fact that the measured values of $w_{rms}$ are identical for the two sets of simulations at different $Re$ demonstrates that viscous effects are negligible, at least for $Fr^{-1}\leq 20$. The right panel shows the results of the suite of experiments at $Pe=0.1$, $Re=600$ (purple symbols), for which $Pr=0.1/600$. We refer to these simulations as ‘diffusive’ runs, because $Pe$ is small. Note that some of these runs were actually integrated using the LPN regime equations instead (square symbols). In that case, the relevant input parameters are $Re$ and $\chi=Pe/Fr^{2}$ ($=BPe$ in the notation of Cope et al.,, 2020). To obtain the corresponding value of $Fr$ for a given $\chi$, simply note that $Fr=\sqrt{Pe/\chi}$ for a given $Pe$. Illustrated as well in the same plots are the various theoretical predictions for the vertical velocity: the red line in each panel corresponds to the SSA theory, while the blue line corresponds to the MSA theory.
@@
Figure 2: Top row: Comparison between the model predictions and the data for non-diffusive simulations with $Pr=0.1$ and two different values of $Re$ (left) and diffusive simulations with $Re=600$, $Pe=0.1$ (right). Symbols show $w_{rms}$ extracted from the DNS and errorbars show the standard deviation of its temporal variability. Squares on the right panel denote LPN simulations (see main text for details). The blue and red lines in each panel show the MSA and SSA scaling predictions, respectively. Bottom row: Regime diagrams for stratified turbulence at $Pr=0.1$ (left) and $Pr=0.1/600\simeq 0.00017$ (right), adapted from Shah et al., (2024). Grey regions support isotropic motions. White regions are viscously controlled ($Re_{b}\leq 1$). Green regions support non-diffusive anisotropic stratified turbulence ($Pe_{b}\geq\mathit{O}(1)$), and purple regions support diffusive anisotropic stratified turbulence ($Pe_{b}\ll\alpha$). The yellow regions are in the ‘intermediate’ regime of Shah et al., (2024) ($\mathit{O}(\alpha)\leq Pe_{b}\ll 1$). Horizontal arrows show the transects through the regimes corresponding to the panels above.
@@
The bottom row of figure 2 shows for comparison the expected regime diagrams for the corresponding values of $Pr$ in each case, based on the asymptotic theory of Shah et al., (2024). The coloured horizontal arrows show the transect taken through parameter space for each series of DNS shown in the top row. The background colours show the expected regime: isotropic motions with $\alpha\simeq 1$ (grey), non-diffusive anisotropic turbulence (green), intermediate regime (yellow), diffusive anisotropic turbulence (violet) and viscous regime (white). The same background colours in the top row show the expected regime transitions as a function of $Fr^{-1}$ at the value of $Pe$ corresponding to the transect taken. We note that while Shah et al., (2024) distinguished the non-diffusive and intermediate regimes, these have the same predicted scalings for $\alpha$ and $w$; in any case, the intermediate regime does not span a large region of parameter space at $Pr=0.1$ and would be difficult to identify even if the scaling laws differed.
@@
3 The effects of intermittency

### v2

@@ Abstract
Recent theoretical progress using multiscale asymptotic analysis has revealed various possible regimes of stratified turbulence. Notably, buoyancy transport can either be dominated by advection or diffusion, depending on the effective Péclet number of the flow. Two types of asymptotic models have been proposed, which yield measurably different predictions for the characteristic vertical velocity and length scale of the turbulent eddies in both diffusive and non-diffusive regimes. The first, termed a ‘single-scale model’, is designed to describe flow structures having large horizontal and small vertical scales, while the second, termed a ‘multiscale model’, additionally incorporates flow features with small horizontal scales, and reduces to the single-scale model in their absence. By comparing predicted vertical velocity scaling laws with direct numerical simulation data, we show that the multiscale model correctly captures the properties of strongly stratified turbulence within regions dominated by small-scale isotropic motions, whose volume fraction decreases as the stratification increases. Meanwhile its single-scale reduction accurately describes the more orderly, layer-like, quiescent flow outside those regions.
keywords:
@@ 1 Introduction
Owing to the associated enhanced rates of irreversible scalar mixing, stratified turbulence is a critical process in the Earth’s atmosphere and oceans, impacting both weather and climate, and in the interiors of stars and gaseous planets, affecting their long-term evolution. Assuming that the buoyancy of the fluid, whether liquid or gaseous, is controlled by a single scalar field, which could be temperature or the concentration of a single solute, the dimensionless Boussinesq equations governing the fluid motions (Spiegel & Veronis, 1960) are
@@
$\boldsymbol{\nabla}\cdot\boldsymbol{u}$ $=0.$ (1c)
@@
Here, $\boldsymbol{u}=(u,v,w)$ is the velocity field expressed in units of $U{\color[rgb]{0,0,0}{}^{*}}$ (where $U{\color[rgb]{0,0,0}{}^{*}}$ is a characteristic horizontal velocity of the large-scale flow), $t$ is the time variable in units of $L{\color[rgb]{0,0,0}{}^{*}}/U{\color[rgb]{0,0,0}{}^{*}}$ (where $L{\color[rgb]{0,0,0}{}^{*}}$ is a characteristic large horizontal scale of the flow), $p$ is the pressure fluctuation away from hydrostatic equilibrium in units of $\rho{\color[rgb]{0,0,0}{}_{m}^{*}}U{\color[rgb]{0,0,0}{}^{*2}}$ (where $\rho{\color[rgb]{0,0,0}{}_{m}^{*}}$ is the mean density of the fluid), and $b$ is the deviation of the buoyancy field away from a linearly stratified background, expressed in units of $L{\color[rgb]{0,0,0}{}^{*}}N{\color[rgb]{0,0,0}{}^{*2}}$ (where $N{\color[rgb]{0,0,0}{}^{*}}$ is the buoyancy frequency of the stable stratification). The $w$ term in the buoyancy equation thus represents the vertical advection of the background stratification. Note that here and throughout this paper, starred quantities are dimensional while non-starred quantities are non-dimensional. The flow is assumed to be driven by a non-dimensional divergence-free horizontal force ${\boldsymbol{F}}_{h}$, which only varies on large spatial scales and long time scales. The unit vector $\boldsymbol{e}_{z}$ points in the direction opposite to gravity. We note that the validity of the Boussinesq approximation in the context of gaseous atmospheric and astrophysical flows (Spiegel & Veronis, 1960) must be verified a posteriori, by checking that the characteristic vertical scale of the flow remains much smaller than a pressure, temperature or density scale height. It is assumed here that $U^{*}$ is always much smaller than the sound speed.
@@
The usual dimensionless governing parameters of the flow emerge; namely, the Reynolds number $Re=U{\color[rgb]{0,0,0}{}^{*}}L{\color[rgb]{0,0,0}{}^{*}}/\nu{\color[rgb]{0,0,0}{}^{*}}$, the Péclet number $Pe=U{\color[rgb]{0,0,0}{}^{*}}L{\color[rgb]{0,0,0}{}^{*}}/\kappa{\color[rgb]{0,0,0}{}^{*}}$ and the Froude number $Fr=U{\color[rgb]{0,0,0}{}^{*}}/N{\color[rgb]{0,0,0}{}^{*}}L{\color[rgb]{0,0,0}{}^{*}}$, where $\nu{\color[rgb]{0,0,0}{}^{*}}$ is the kinematic viscosity, and $\kappa{\color[rgb]{0,0,0}{}^{*}}$ is the buoyancy diffusivity, while the Prandtl number, $Pr=\nu{\color[rgb]{0,0,0}{}^{*}}/\kappa{\color[rgb]{0,0,0}{}^{*}}=Pe/Re$ is a property of the fluid. Typically, $Pr\sim O(1)$ in air and water, but is very small in astrophysical fluids (of order $10^{-2}$ in degenerate plasmas and liquid metals, and much smaller in non-degenerate stellar plasmas, see Lignières, 2020).
@@
As the stratification increases ($Fr\rightarrow 0$), vertical motions are increasingly suppressed and restricted to small characteristic vertical scales $l_{z}=O(\alpha)$, where the emergent aspect-ratio $\alpha{\color[rgb]{0,0,0}=l_{z}^{*}/L^{*}}$ is an increasing function of $Fr$ but could also depend on $Re$ and $Pe$. In the limit $(\alpha,Fr)\rightarrow 0$, asymptotic analysis has successfully been used to derive reduced equations for stratified turbulence and to gain insight into its properties. Brethouwer et al. (2007), following Billant & Chomaz (2001), proposed an asymptotic reduction in which the vertical coordinate $z$ is rescaled as $\zeta=z/\alpha$ (with $\alpha\ll 1$), while the horizontal coordinates ${\boldsymbol{x}}_{h}=(x,y)$ remain of order unity. Accordingly, all dependent variables $q$ are expressed as $q({\boldsymbol{x}}_{h},\zeta,t;\alpha)$. As in Klein (2010), we refer to this type of model, which is designed to capture the essence of a scale-specific process, as a single-scale asymptotic (SSA) model. This vertical rescaling, when used in (1), reveals the importance of the emergent buoyancy Reynolds and Péclet numbers, defined here as
@@
respectively. Brethouwer et al. (2007) showed that balancing the mass continuity equation in the limit $\alpha\rightarrow 0$ requires $w=O(\alpha)$. When $Re_{b}$ and $Pe_{b}$ are at least $O(1)$, dominant balance in the buoyancy equation implies $b=O(\alpha)$. Finally, the vertical component of the momentum equation reduces to hydrostatic equilibrium when $\alpha\rightarrow 0$, yielding $\alpha=Fr$, as first argued in the inviscid and non-diffusive case by Billant & Chomaz (2001).
More recently, Shah et al. (2024) noted that in the limit of $Pr\ll 1$, it is possible to have a regime in which $Pe_{b}\ll 1\leq Re_{b}$. They demonstrated that in this case, the SSA model and corresponding asymptotic expansion reveal instead that $w=O(\alpha)$ and $b=O(\alpha Pe_{b})$, with $\alpha=(Fr^{2}/Pe)^{1/4}$ (see also Lignières, 2020; Skoutnev, 2023).
Crucial to the SSA theory is the notion that every component of the flow is strongly anisotropic, with large horizontal scales and a small vertical scale. In this model, therefore, the vertical fluid motions are primarily driven by the divergence of the horizontal flow, as illustrated schematically in figure 1. Chini et al. (2022), however, noted that the SSA theory ignores the possibility that isotropic motions with small horizontal scales may also exist and, in fact, are commonly seen in numerical simulations of stratified turbulence at sufficiently large $Re_{b}$ (cf. Maffioli & Davidson, 2016; Cope et al., 2020; Garaud, 2020). They proposed a new asymptotic reduction that explicitly incorporates two horizontal scales and two time scales, such that all dependent variables are expressed as
@@
where the subscripts $s$ and $f$ are used to denote slow and fast scales, respectively. Again following Klein (2010), we refer to the resulting reduced equations as a multiscale asymptotic (MSA) model. Chini et al. (2022) showed that these definitions imply that the large-scale motions remain strongly anisotropic with an aspect ratio $\alpha$, as in the SSA model, but can coexist with isotropic small-scale motions that evolve on the fast time scale $t_{f}$, and vary on the small vertical coordinate $\zeta$ and the ‘fast’ horizontal coordinate ${\boldsymbol{x}}_{f}$. These small-scale motions are self-consistently driven by an instability of the local vertical shear emergent from the larger-scale horizontal flow (see figure 1), and are gradually stabilized as the stratification increases at fixed $Re$. They essentially disappear beyond a certain threshold, at which point the MSA model naturally recovers the SSA model and its predicted scalings. For the sake of clarity, however, we refer in what follows to the SSA model and its scalings whenever small horizontal scales are dynamically negligible, and to the MSA model and its scalings whenever they are dynamically important, even though the MSA model does in fact naturally cover both cases.
@@
In the asymptotic limit where $Re_{b}\geq O(1)$ and $Pe_{b}\geq O(1)$, Chini et al. (2022) found that $w=O(\alpha^{1/2})$, $b=O(\alpha)$ and $\alpha=Fr$ when $\alpha\rightarrow 0$. Their scaling prediction for $w$ thus deviates substantially from that of Brethouwer et al. (2007), but recovers that of Riley & Lindborg (2012) albeit using different arguments (for details see Shah et al., 2024). That scaling has been tentatively validated by Maffioli & Davidson (2016) in run-down direct numerical simulations (DNS) of stratified turbulence.
@@
Extending the MSA theory to the low $Pr$ case, Shah et al. (2024) recovered the results of Chini et al. (2022) when $Pe_{b}\geq\mathit{O}(1)$. They also found that $\alpha=Fr$ and $w=O(\alpha^{1/2})$ both continue to hold in an ‘intermediate’ regime where $O(\alpha)\leq Pe_{b}\ll 1$. However, when $Pe_{b}\ll\alpha$, a new fully diffusive regime emerges in which $w=O(\alpha^{1/2})$, $b=O(Pe_{b}\alpha^{1/2})$ and $\alpha=(Fr^{2}/Pe)^{1/3}$. As for the $Pr=\mathit{O}(1)$ scenario, the scaling predictions of the MSA theory differ substantially from those emerging from the low $Pe_{b}$ limit of the SSA theory but recover them when small scales are absent. The various theories and their predicted scalings for $Re_{b}\geq O(1)$, with $Pe_{b}\geq O(\alpha)$ or $Pe_{b}\ll\alpha$, respectively, are summarized in figure 1.
@@
Therefore, an interesting question is whether evidence for these scaling laws can be found in DNS data. Recently, two series of DNS were presented by Cope et al. (2020) and Garaud (2020), respectively, which solved equations (1) with $\boldsymbol{F}_{h}=\sin(y)\boldsymbol{e}_{x}$ (where $\boldsymbol{e}_{x}$ is a unit vector in the streamwise, i.e. $x$, direction; see §2 for further details). In their $Pe<1$ simulations (where by construction $Pe_{b}\ll\alpha$), Cope et al. (2020) found that the vertical length scale of the turbulent motions scales as $\alpha=(Fr^{2}/Pe)^{1/3}$, validating the predictions of Shah et al. (2024) in that limit. This scaling, however, was not as clearly evident in the high $Pe$ but low $Pe_{b}$ data of Garaud (2020). One potential explanation is that $Re_{b}$ is relatively low in these simulations (which have $Pr=0.1$, so $Re_{b}=10Pe_{b}$), implying viscous effects are not necessarily negligible. In the limit of high $Pe_{b}$, Garaud (2020) was unable to find evidence for the $\alpha=Fr$, $w=O(Fr)$ scaling of Brethouwer et al. (2007) and was unaware at the time of the scaling $w=O(Fr^{1/2})$ obtained by Chini et al. (2022), proposing instead on empirical grounds that $w\propto\alpha=Fr^{2/3}$ provides the best fit to the data. The apparent discrepancy between Garaud’s data and previous models therefore prompts us to analyze some new DNS results and to revisit the available data from Cope et al. (2020) and Garaud (2020) in the light of the MSA models of stratified turbulence recently derived by Chini et al. (2022) and Shah et al. (2024).
@@ 2 Comparison of theory with DNS
In what follows, we extend and re-analyze the datasets presented in Cope et al. (2020) and Garaud (2020). Both studies performed DNS of the set of non-dimensional equations (1) with $\boldsymbol{F}_{h}=\sin(y)\boldsymbol{e}_{x}$ in a triply-periodic domain of size $L_{x}=4\pi,L_{y}=2\pi,L_{z}=2\pi$ using the PADDI code (Traxler et al., 2011). With this choice, $L^{*}$ is simply the inverse horizontal wavenumber $k^{*}_{f}$ of the forcing, and the dimensional velocity scale $U^{*}$ is $(F_{0}^{*}/\rho^{*}_{m}k^{*}_{f})^{1/2}$, where $F^{*}_{0}$ is the dimensional amplitude of the forcing. In both studies, the streamwise length of the domain $L_{x}=4\pi$ was chosen to be close to the period of the fastest-growing mode of the horizontal shear instability of the Kolmogorov flow driven by that body force (see e.g. Cope et al., 2020). This ensures the natural generation of large-scale flows in the horizontal direction, whose nonlinear evolution then generates flows on $O(1)$ scales in both the $x$ and $y$ directions. As illustrated in the Appendix, we have verified that the results in the non-viscous, non-diffusive regime are essentially independent of $L_{x}$ and $L_{y}$ as long as the domain is large enough to allow that primary mode of horizontal shear instability to grow.
[…]
@@
To complement this dataset, we have run additional simulations at $Re=1000$, $Pe=100$. These DNS have twice the spatial resolution of those of Cope et al. (2020) and Garaud (2020) and, thus, have only been integrated for up to 50 time units in the statistically stationary regime. In addition, the full fields are too large to be saved regularly, so we have saved two-dimensional slices through the data in the $(x,y)$, $(y,z)$ and $(x,z)$ planes. These simulations are only used for visualizations and to assess the influence of viscosity by comparison with the $Re=600$ results. Note that for these new $Re=1000$ runs, and for all previously published ones in Cope et al. (2020) and Garaud (2020), we have ensured that the product of the maximum resolved wavenumber and the Kolmogorov scale is always greater than one, ensuring that the flow field (and the buoyancy field, since $Pr\leq 0.1$ in all cases), is resolved down to the dissipation scales (see Cope et al., 2020; Garaud, 2020, for details).
@@
We compare the $w_{rms}$ data and the various theories in the top row of figure 2. The left panel shows $w_{rms}$ as a function of $Fr^{-1}$, measured for the $Pe=60$, $Re=600$ runs (green symbols), and for the $Pe=100$, $Re=1000$ runs (orange symbols). Note that $Pr=0.1$ in both cases. We refer to these simulations as ‘non-diffusive’ because $Pe$ is large. The fact that the measured values of $w_{rms}$ are identical for the two sets of simulations at different $Re$ demonstrates that viscous effects are negligible, at least for $Fr^{-1}\leq 20$. The right panel shows the results of the suite of experiments at $Pe=0.1$, $Re=600$ (purple symbols), for which $Pr=0.1/600$. We refer to these simulations as ‘diffusive’ runs, because $Pe$ is small. Note that some of these runs were actually integrated using the LPN regime equations instead (square symbols). In that case, the relevant input parameters are $Re$ and $\chi=Pe/Fr^{2}$ ($=BPe$ in the notation of Cope et al., 2020). To obtain the corresponding value of $Fr$ for a given $\chi$, simply note that $Fr=\sqrt{Pe/\chi}$ for a given $Pe$. Illustrated as well in the same plots are the various theoretical predictions for the vertical velocity: the red line in each panel corresponds to the SSA theory, while the blue line corresponds to the MSA theory.
@@
Figure 2: Top row: Comparison between the model predictions and the data for non-diffusive simulations with $Pr=0.1$ and two different values of $Re$ (left) and diffusive simulations with $Re=600$, $Pe=0.1$ (right). Symbols show $w_{rms}$ extracted from the DNS and errorbars show the standard deviation of its temporal variability. Squares on the right panel denote LPN simulations (see main text for details). The blue and red lines in each panel show the MSA and SSA scaling predictions, respectively. Bottom row: Regime diagrams for stratified turbulence at $Pr=0.1$ (left) and $Pr=0.1/600\simeq 0.00017$ (right), adapted from Shah et al. (2024). Grey regions support isotropic motions. White regions are viscously controlled ($Re_{b}\leq 1$). Green regions support non-diffusive anisotropic stratified turbulence ($Pe_{b}\geq\mathit{O}(1)$), and purple regions support diffusive anisotropic stratified turbulence ($Pe_{b}\ll\alpha$). The yellow regions are in the ‘intermediate’ regime of Shah et al. (2024) ($\mathit{O}(\alpha)\leq Pe_{b}\ll 1$). Horizontal arrows show the transects through the regimes corresponding to the panels above.
@@
The bottom row of figure 2 shows for comparison the expected regime diagrams for the corresponding values of $Pr$ in each case, based on the asymptotic theory of Shah et al. (2024). The coloured horizontal arrows show the transect taken through parameter space for each series of DNS shown in the top row. The background colours show the expected regime: isotropic motions with $\alpha\simeq 1$ (grey), non-diffusive anisotropic turbulence (green), intermediate regime (yellow), diffusive anisotropic turbulence (violet) and viscous regime (white). The same background colours in the top row show the expected regime transitions as a function of $Fr^{-1}$ at the value of $Pe$ corresponding to the transect taken. We note that while Shah et al. (2024) distinguished the non-diffusive and intermediate regimes, these have the same predicted scalings for $\alpha$ and $w$; in any case, the intermediate regime does not span a large region of parameter space at $Pr=0.1$ and would be difficult to identify even if the scaling laws differed.
@@
3 Turbulent patches vs. quiescent flow